# Entity or Relation Embeddings? An Analysis of Encoding Strategies for Relation Extraction

### v1

@@ 4 Experiments
In this section, we compare the effectiveness of the considered relation embedding strategies on a number of standard relation extraction benchmarks. Our main hypothesis is that the common [H,T] strategy essentially leads models to focus on the semantic types of the head and tail entity. We are thus interested in comparing the [H,T] and EnCore strategies. Furthermore, we hypothesise that the MASK strategy leads to embeddings capturing information that is complementary to that captured by the [H,T] embeddings. Accordingly, we are interested in comparing [H,T] with [H,T,Mask] and [H,T]+Mask. We first present the details of our experimental setup in Section 4.1. Section 4.2 subsequently summarises our main results. Finally, we present some additional analysis in Section 4.3.
@@ 4.1 Experimental Setup
We consider a number of standard benchmarks for relation extraction. First, we use TACRED Zhang et al. (2017), a popular relation extraction benchmark. Two revisions of this dataset have been proposed, both of which are aimed at addressing some of the issues that arose in the original dataset due to the use of crowd-sourcing. First, TACREV Alt et al. (2020) was obtained by re-annotating the 5000 most challenging instances from the development and test sets. Specifically, to select these instances, the authors looked at how often models disagreed with each other and with the ground truth. Re-TACRED Stoica et al. (2021) was obtained by re-annotating the entire dataset. Following the tradition in recent work on relation extraction, we report results on all three variants in terms of F1 score. Next, we evaluate on a distantly supervised dataset that was introduced by Sorokin and Gurevych (2017) by aligning Wikipedia and Wikidata (Wiki-WD). Finally, we will also consider the distantly supervised dataset that was introduced by Riedel et al. (2010) by aligning articles from the New York Times with Freebase (NYT-FB). Following the tradition from previous work, we report the results on Wiki-WD in terms of F1 score and the results on NYT-FB in terms of precision at 10 (P@10) and 30 (P@30), averaged across all relation labels. Table 1 summarises the main characteristics of the considered datasets.
@@
FOR TACRED and its variants, we consider a number of recent baselines. First, we include a comparison with SpanBERT Joshi et al. (2020) and LUKE Yamada et al. (2020). Next, we consider the Typed Marker strategy from Zhou and Chen (2022), and the Curriculum Learning variant from Park and Kim (2021). Note, however, that these last two methods are not directly comparable with our methods, as they rely on the gold entity type labels that are provided as part of the TACRED dataset. We do not use these labels for our models since such information is typically not available in practice. For Wiki-WD and NYT-FB, we compare with RECON Bastos et al. (2021) and KGPool Nadgeri et al. (2021). Note, however, that these state-of-the-art methods are again not directly comparable with our methods. They are focused on modelling and exploiting knowledge from the Wikidata and Freebase knowledge graphs. We do not take such information into account, as our focus is on comparing different encoding strategies for learning relation embeddings.
Training Details
Unless stated otherwise, the relation encoder is first pre-trained using the strategy from Section 3.1, before being fine-tuned, as explained in Section 3.2. In our default setting, we use the same training set for pre-training and fine-tuning the relation encoder. For TACRED, TACREV and ReTACRED, where the training set is comparatively smaller, we also experiment with a variant where we instead use the Gigaword corpus for pre-training the relation encoder, following the self-supervision strategy from Section 3.1. Finally, we also report results where the pre-training step is omitted and we directly train the relation encoder using the fine-tuning strategy from Section 3.2.
For our main experiments, we use roberta-large [^4] https://huggingface.co/docs/transformers/model_doc/roberta to initialise the relation encoder. We pre-train the model for 25 epochs and select the checkpoint with the minimum validation loss. For the fine-tuning step, we similarly train the model for 25 epochs and select the best checkpoint based on the validation set. We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $45e-$ 4 and a weight decay $\lambda$ of $51e-$ 5. The temperature $\tau$ in the contrastive loss was set to 0.05. For the relation classifier, we used 256-dimensional embeddings for the hidden layer.
@@ 4.2 Results
SpanBERT 70.8 $*{}^{*}$ ${}^{\dagger}$
LUKE 72.7 ${}^{\circ}$ ${}^{\circ}$
@@
Mask 22.8 22.4 22.9
[H,T] 60.7 62.9 71.9
[H,T,Mask] 72.7 72.8 81.3
[H,T]+Mask 77.6 83.9 91.5
EnCore+Mask 77.9 83.5 91.6
EnCore+[H,T]+Mask 77.8 83.3 91.5
@@
Mask 24.0 23.1 37.8
[H,T] 62.9 65.9 79.0
[H,T,Mask] 74.1 74.7 82.6
[H,T]+Mask 78.1 83.3 92.4
EnCore+Mask 78.9 84.4 93.1
EnCore+[H,T]+Mask 78.9 84.1 92.7
@@
Mask 49.8 48.2 48.4 51.9 49.3
[H,T] 75.7 74.2 74.9 79.3 78.0
[H,T,Mask] 81.8 80.3 81.1 87.9 85.8
[H,T]+Mask 89.3 88.9 89.1 94.2 92.9
EnCore+Mask 89.0 88.6 88.8 94.8 92.9
EnCore+[H,T]+Mask 89.4 88.0 88.7 94.3 92.6
@@
Table 2 summarises the results for TACRED, TACREV and Re-TACRED. Let us first consider the standard pre-training results (i.e. where the models are pre-trained on the standard training set). A number of observations stand out. First, the Mask strategy on its own performs poorly. Second, there is clear evidence that the [H,T] and Mask strategies are complementary: both [H,T,Mask] and [H,T]+Mask substantially outperform [H,T] and Mask on their own. We can furthermore see that [H,T]+Mask outperforms [H,T,Mask]. The combined pre-training strategy which is used by [H,T,Mask] thus indeed seems to largely fail to learn meaningful [MASK] embeddings. Finally, the EnCore+Mask strategy matches or even outperforms [H,T]+Mask. This is surprising, given that the EnCore embeddings are specifically trained to capture semantic types and are not fine-tuned on the relation extraction datasets. The strong performance of EnCore+Mask thus clearly supports the idea that the [H,T] representation mostly captures the entity types, rather than the relationship itself.
@@
The results of the configurations that were pre-trained on Gigaword show the effectiveness of this pre-training strategy, as the results of all models are better than with standard pre-training. While the fact that pre-training on external datasets can bring benefits is in itself unsurprising, this self-supervision strategy based on coreference chains was not previously tested for relation extraction. It offers a convenient way to improve relation extraction systems, since it does not rely on an entity linked corpus. Comparing the performance of the different configurations, after pre-training on Gigaword, we see the same patterns as with standard pre-training. The Encore+Mask strategy emerges as the best model overall, which in particular confirms the usefulness of combining entity typing information with direct relation embeddings.
@@
We can also see that forgoing pre-training altogether has a detrimental effect. A direct comparison with the baselines is difficult, as the two strongest baselines use additional information (i.e. the gold entity type labels). Nonetheless, we can see that our best configurations outperform the state-of-the-art methods. The improvements are relatively small for TACREV and Re-TACRED, especially with standard pre-training, while being somewhat larger for TACRED.
@@
4.3 Analysis

### v2

@@ 4 Experiments
We compare the effectiveness of the considered relation embedding strategies on a number of standard relation extraction benchmarks. Our main hypothesis is that the common [H,T] strategy essentially leads models to focus on the semantic types of the head and tail entity. We are thus interested in comparing the [H,T] and EnCore strategies. Furthermore, we hypothesise that the information captured by the Mask embeddings is complementary to that captured by the [H,T] embeddings. Accordingly, we are interested in comparing [H,T] with [H,T,Mask] and [H,T]+Mask.
@@ 4.1 Experimental Setup
We evaluate on five standard benchmarks. First, we use TACRED Zhang et al. (2017), a popular relation extraction benchmark. Two revisions of this dataset have been proposed, both of which are aimed at addressing problems with noisy annotations. In particular, TACREV Alt et al. (2020) was obtained by re-annotating the 5000 most challenging instances from the development and test sets. Specifically, to select these instances, the authors looked at how often models disagreed with each other and with the ground truth. Re-TACRED Stoica et al. (2021) was obtained by re-annotating the entire dataset. Following tradition, we report results on all three variants in terms of F1 score. Next, we evaluate on a distantly supervised dataset that was introduced by Sorokin and Gurevych (2017) by aligning Wikipedia and Wikidata (Wiki-WD). Finally, we also consider the distantly supervised dataset that was introduced by Riedel et al. (2010) by aligning articles from the New York Times with Freebase (NYT-FB). The NYT-FB dataset does not have an explicit validation set. As a result, we keep 10% of the training set as a validation set and train on the remaining 90%. Following the tradition from previous work, we report the results on Wiki-WD in terms of F1 score and the results on NYT-FB in terms of precision at 10 (P@10) and 30 (P@30), averaged across all relation labels. Table 1 summarises the main characteristics of the considered datasets.
@@
FOR TACRED and its variants, we consider a number of recent baselines. First, we include a comparison with SpanBERT Joshi et al. (2020) and LUKE Yamada et al. (2020). We furthermore compare with KnowBERT Chen et al. (2022), which also uses a prompt with the [MASK] token. They improve on the standard Mask strategy, among others, by incorporating predicted entity types, and are thus a natural baseline for our methods. Finally, we consider the Typed Marker strategy from Zhou and Chen (2022), and the Curriculum Learning variant from Park and Kim (2021). Note, however, that these last two methods are not directly comparable with our methods, as they rely on the gold entity type labels that are provided as part of the TACRED dataset. We do not use these labels for our models since such information is typically not available in practice. For Wiki-WD and NYT-FB, we compare with RECON Bastos et al. (2021) and KGPool Nadgeri et al. (2021). Note, however, that these state-of-the-art methods are again not directly comparable with our methods. They are focused on modelling and exploiting knowledge from the Wikidata and Freebase knowledge graphs. We do not take such information into account, as our focus is on comparing different encoding strategies for learning relation embeddings.
Training Strategies
Unless stated otherwise, the relation encoder is first pre-trained using the strategy from Section 3.1, before being fine-tuned, as explained in Section 3.2. In our default setting, we use the same training set for pre-training and fine-tuning the relation encoder. For TACRED, TACREV and ReTACRED, where the training set is comparatively smaller, we also experiment with a variant where we instead use the Gigaword corpus for pre-training the relation encoder, following the self-supervision strategy from Section 3.1. Finally, we also report results where the pre-training step is omitted and we directly train the relation encoder using the fine-tuning strategy from Section 3.2. For our main experiments, we use roberta-large [^5] https://huggingface.co/docs/transformers/model_doc/roberta to initialise the relation encoder.
@@ 4.2 Results
SpanBERT 70.8 78.0 85.3
LUKE 72.7 80.6 90.3
KnowPrompt 72.4 81.4 90.9
@@
Mask 23.3 22.9 23.2
[H,T] 61.4 63.6 72.3
[H,T,Mask] 73.0 73.8 81.9
[H,T]+Mask 78.5 84.6 91.9
EnCore+Mask 78.9 83.9 91.8
EnCore+[H,T]+Mask 78.1 84.1 91.8
@@
Mask 24.6 23.5 38.2
[H,T] 63.2 66.3 79.7
[H,T,Mask] 74.4 75.1 82.9
[H,T]+Mask 78.5 83.6 92.7
EnCore+Mask 79.1 84.9 93.5
EnCore+[H,T]+Mask 79.0 84.4 93.2
@@
Mask 50.3 48.7 48.9 52.3 49.7
[H,T] 75.9 74.7 75.2 79.6 78.5
[H,T,Mask] 82.1 80.8 81.7 88.4 86.1
[H,T]+Mask 89.8 89.3 89.6 94.7 93.1
EnCore+Mask 89.4 88.8 89.1 94.9 93.0
EnCore+[H,T]+Mask 89.8 88.6 88.9 94.7 92.9
@@
Sentence Label
Asia Bibi was sentenced to hang in Pakistan’s central province of Punjab earlier this month after being accused of insulting the Prophet Mohammed in 2009. Gold: no relation
Mask: per:origin
[H,T]: per:state or provinces of residence
[H,T]+Mask: no relation
In October, she filed a complaint with the police in Rio saying he had kidnapped her and tried to threaten her into having an abortion. Gold: no relation
Mask: per:origin
[H,T]: per:cities of residence
[H,T]+Mask: no relation
Benjamin Chertoff is the Editor in Chief of Popular Mechanics magazine as well as the cousin of the Director of Homeland Security, Michael Chertoff. Gold: no relation
Mask: per:employee of
[H,T]: per:employee of
[H,T]+Mask: no relation
WASHINGTON – The National Restaurant Association gave \$35,000 – a year’s salary – in severance pay to a female staff member in the late 1990s after an encounter with Herman Cain, its chief executive at the time, made her uncomfortable working there, three people with direct knowledge of the payment said on Tuesday. Gold: org:top members/employees
Mask: per:employee of
[H,T]: no relation
[H,T]+Mask: org:top members/employees
“From January 1, I, Charles Ble Goude and the youth of Ivory Coast are going to liberate the Golf Hotel with our bare hands," the political showman turned minister declared Wednesday, to a cheering crowd of hardline supporters. Gold: per:title
Mask: per:employee of
[H,T]: no relation
[H,T]+Mask: per:title
Table 4: Comparison of the TACREV test set predictions from the Mask, [H, T] and [H, T] + Mask models that were initialised using RoBERTa-large.
@@
Table 2 summarises the results for TACRED, TACREV and Re-TACRED. Let us first consider the standard pre-training results (i.e. where the models are pre-trained on the standard training set). A number of observations stand out. First, the Mask strategy on its own performs poorly. Second, there is clear evidence that the [H,T] and Mask strategies are complementary: both [H,T,Mask] and [H,T]+Mask substantially outperform [H,T] and Mask on their own. We can furthermore see that [H,T]+Mask outperforms [H,T,Mask]. The combined pre-training strategy which is used by [H,T,Mask] thus indeed seems to largely fail to learn meaningful [MASK] embeddings. Finally, the EnCore+Mask strategy matches or even outperforms [H,T]+Mask. This is surprising, given that the EnCore embeddings are specifically trained to capture semantic types and are not fine-tuned on the relation extraction datasets. The strong performance of EnCore+Mask thus clearly supports the idea that the [H,T] representations mostly capture the entity types, rather than the relationship itself. We can similarly see that EnCore+[H,T]+Mask does not generally improve on Encore+Mask, which further supports this idea.
@@
Pre-training on Gigaword leads to clear and consistent improvements, compared to standard pre-training. While the fact that pre-training on external datasets can bring benefits is in itself unsurprising, this self-supervision strategy based on coreference chains was not previously tested for relation extraction. It offers a convenient way to improve relation extraction systems, since it does not rely on an entity linked corpus. Comparing the performance of the different configurations, after pre-training on Gigaword, we see the same patterns as with standard pre-training. The Encore+Mask strategy emerges as the best model overall, which in particular confirms the usefulness of combining entity embeddings information with relation embeddings.
@@
We can also see that forgoing pre-training altogether has a detrimental effect. A direct comparison with the baselines is difficult, as the two strongest baselines use additional information (i.e. the gold entity type labels). Nonetheless, our best configurations consistently outperform the state-of-the-art methods, with the improvements being most pronounced for TACRED.
@@
Further analysis of our results can be found in the appendix. Among others, we show that the conclusions from this section remain valid when other language models than RoBERTa-large are used as the encoder. We also present a detailed error analysis to support our claims about the limitations of the [H,T] and Mask strategies, an analysis of the models in a setting with limited training data, and an analysis of the impact of the dimensionality of the entity and relation embeddings.
4.3 Qualitative Analysis
In Table 4, we compare predictions of the Mask, [H,T] and [H,T]+Mask models for the TACREV test set. The first three cases illustrate how the [H,T] model often overly relies on the semantic types of the entities, without fully taking into account the actual sentence context, a problem known as entity bias Wang et al. (2022b). In particular, as the first two examples illustrate, given a person and a place, the [H,T] model frequently predicts that the individual is a resident of that place, even if there is no relationship expressed in the given context. In contrast, [H,T]+Mask correctly predicts no relation in these cases. The third example shows a similar issue, which arises when the sentence refers to a person and an organisation. In this case, the [H,T] model incorrectly predicts that the person is employed by that organisation. As the fourth example illustrates, however, the opposite situation also arises, where the [H,T] model incorrectly predicts no relation. Furthermore, several examples illustrate how the Mask model struggles because it does not adequately capture the semantic types of the entities. For instance, in the first two examples, the model predicts per:origin despite the fact that the tail entity is not a country. The issue is most clearly illustrated by the fifth example, where the Mask model predicts employee of, despite the tail entity not being an organisation. Overall, these examples support the view that the [H,T] model focuses too much on modelling the entity types, which is not always sufficient, while conversely, Mask struggles because it does not sufficiently take entity types into account. We include further examples in the appendix, which further support our findings.
5 Conclusions
Limitations
Our analysis in this paper was limited to the English language, and we only considered the setting of fully supervised sentence-level relation extraction. Since our focus was on learning representations (i.e. relation embeddings), it is not straightforward to transfer our findings to the zero-shot relation extraction setting (as we do not attempt to model the relation labels). We did not consider the use of LLMs for relation extraction in this paper. On the one hand, this is due to the fact that applying LLMs to this setting is not straightforward, especially when the set of candidate relation labels is large. Progress in this area has indeed been slow, as we highlighted in the related work section. Moreover, while LLMs can be used to extract symbolic representations (e.g. knowledge graph triples), they are often less suitable for learning embeddings. Traditionally, relation embeddings have primarily been used as an intermediate representations, before relation labels were predicted, and from this perspective, we may wonder whether relation embeddings are still needed in the LLM era. However, beyond acting as an intermediate representation, embeddings have a number of important advantages. They can, in principle, capture much more subtle distinctions than is possible with pre-defined discrete relation labels. As such, they are more suitable for modelling relational similarity (i.e. analogy), for instance.
Acknowledgements
This research was supported by EPSRC grant EP/W003309/1 and undertaken using the supercomputing facilities at Cardiff University operated by Advanced Research Computing at Cardiff (ARCCA) on behalf of the Cardiff Supercomputing Facility and the HPC Wales and Supercomputing Wales (SCW) projects. We acknowledge the support of the latter, which is part-funded by the European Regional Development Fund (ERDF) via the Welsh Government.
Appendix A Training Details
We pre-train the model for 25 epochs and select the checkpoint with the minimum validation loss. For the fine-tuning step, we similarly train the model for 25 epochs and select the best checkpoint based on the validation set. We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $45e-$ 4 and a weight decay $\lambda$ of $51e-$ 5. The temperature $\tau$ in the contrastive loss was set to 0.05.
Appendix B Analysis
In this section we provide some further analysis of the different relation embeddings. In particular, we make the following observations:
  • By analysing model outputs, we find that the “no relation” label (from TACRED) is the most difficult label for all variants, where models are particularly prone to predicting some relationship even if none is expressed in the sentence. The [H,T]+Mask variant suffers considerably less from this problem, which helps to explain its outperformance over [H,T].
  • In a qualitative analysis of the model outputs, we provide further evidence that [H,T] focuses too much on the entity types. In particular, the model frequently predicts a relationship that holds for entities of the same type but is not expressed in the sentence.
  • We show that the predictions of the masked language model for the [MASK] token are largely meaningless, even though the contextualised representation of this token is useful for relation classification. This is expected, given that the model is not trained to produce meaningful verbalisations of the relationship.
Confusion Matrices
In Figures 34 and 5, we show the confusion matrix for 5 randomly sampled relations and the “no relation” label, for the TACREV test set. For this analysis, we focus on the Mask, [H,T] and [H,T]+Mask models. We refer to any label predictions that fall outside of the six sampled classes as the “other" class. We can clearly observe that the “no relation" label is the most difficult label for all three models: most errors occur when the models predict a relationship while none exists, or vice versa. We can also see that the [H,T]+Mask model regularly outperforms [H,T], which in turn outperforms the Mask model when it comes to modelling the “no relation" label.
Error Comparison
In Figure 6, we present the analysis of model performance that highlights instances where one model outperforms the other. Specifically, we focus on the [H,T] and [H,T]+mask models, showing how often one model makes a mistake while the other gets it right, for the TACREV test set. Compatible with our findings from the confusion matrices, the no “relation” label proves challenging for both methods. There are several instances where [H,T] makes a mistake while [H,T]+Mask does not, and vice versa, although [H,T]+Mask overall performs better. When we look at the other relation labels, however, we can see a clear pattern, as there are only very few instances where [H,T]+Mask makes a mistake without [H,T] also making the same mistake. This shows that the improvement of [H,T]+Mask over [H,T] is highly consistent.
Qualitative Analysis of Model Outputs
Overall, we find that around 70% of the errors of the [H,T] model are “no relation” misclassifications, where most of the remaining mistakes arise because the model confuses relations between entities of the same type, including:
  • family relationships such as “per:siblings” and “per:children”;
  • relationships linking people to geographic regions, such as “per:state or province of death” and “per:state or province of residence”;
  • relationships linking people to organisations, such as “org:founded by” and “org:top members/employees”.
We provide several examples of both types of errors. In particular, Table 5 shows cases where [H,T] mistakenly predicts the “no relation” label, while Table 6 focuses on examples where the [H,T] model confuses the target relation with a relation between entities of the same type.
The Mask model performs worse overall, and there is less of a pattern in the types of errors it makes. Similar to the other models, it is also prone to incorrectly predicting “no relation”. Several examples of this can be seen in Tables 5 and 6. One particular weakness of Mask is that it sometimes fails to correctly predict the direction of a relationship, confusing the target relation with its inverse. This can be seen, for instance, in the last example of Table 6, where Mask confuses “per:children” with “per:parents”. Another example can be found in Table 4 (fourth instance), where the Mask model confused “org:top members/employees” with “per:employee of”. As we highlighted in our qualitative analysis in the main paper, Mask also makes mistakes because it fails to take into account the semantic types of the entities.
Encoder Predictions for the [MASK] token
In Table 7, we present examples of the tokens which are predicted by the language model for the [MASK] position of the appended relation prompts. Specifically, we use our pre-trained [H, T] + Mask encoder that was initialised using RoBERTa-large. We consider the top five tokens directly predicted for the [MASK] positions by the encoder. We compare these predictions with the predictions from our full [H, T] + Mask model, which comes with a classifier on top of the encoder. As can be seen, the token predictions do not adequately capture the relationships that are expressed in the given sentences. This is to be expected, since the InfoNCE loss which is used during pre-training encourages the embeddings of sentences that express the same relationship to be similar, but these vectors are no longer aligned with the tokens from the encoder’s vocabulary. This illustrates the need for a classifier that maps the embeddings of the [MASK] token in the relation prompt to dataset-specific relation labels.
Figure 3: The confusion matrix for the Mask model on the TACREV test set.
Figure 6: The comparison of the errors made by the [H,T] and [H,T]+Mask models on the TACREV test set.
Sentence Label
“The current attempt to restore the commission was masterminded by a suspected mobster, Matteo Messina Denaro, who is among a handful of people vying to replace Provenzano," police said. Gold: per:title
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:title
She ended up leaving Iraq under the threat of losing her job and returning home to Texas to seek medical and psychiatric treatment for post traumatic stress syndrome. Gold: per:state or provinces of residence
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:state or provinces of residence
Wen was tried with his wife, Zhou Xiaoya, and three former Chongqing police associates all of whom received jail sentences of up to 20 years. Gold: per:spouse
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:spouse
Knox’s father, Curt Knox, said his daughter looked “confident in what she wants to say." Gold: per:children
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:children
Her mother, 60-year-old Claudie Mamane, tried to jump from the van while it was still moving and injured her arm. Gold: per:parents
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:parents
He is also survived by his parents and a sister, Karen Lange, of Washington, and a brother, Adam Lange, of St. Louis. Gold: per:siblings
Mask: per:parents
[H,T]: no relation
[H,T]+Mask: per:siblings
After more than 20 years of wearing the same long hairstyle, silver-tressed Ponce Kiah Marchelle Heloise Cruse Evans, known simply as Heloise, has a new ’do for the New Year. Gold: per:alternate names
Mask: no relation
[H,T]: no relation
[H,T]+Mask: per:alternate names
Table 5: Comparison of the TACREV test set predictions from the Mask, [H, T] and [H, T] + Mask models that were initialised using RoBERTa-large, focusing on examples where [H,T] incorrectly predicts “no relation”.
Sentence Label
A man who shot and killed three women in a Pennsylvania health club, then himself, apparently blogged his rage-filled preparations –with the final chilling entry announcing the “big day" of the massacre. Gold: per:state or province of death
Mask: no relation
[H,T]: per:state or provinces of residence
[H,T]+Mask: per:state or province of death
Her brother-in-law, Wen, served as a top Chongqing police official for 16 years before taking over the city’s judiciary. Gold: per:other family
Mask: per:siblings
[H,T]: per:siblings
[H,T]+Mask: per:other family
Robert Holden, deputy director at the National Congress of American Indians, said “the Washington DC-based group is hopeful that the use of secured cards could be expanded to allow tribal members to travel abroad." Gold: org:state or province of headquarters
Mask: no relation
[H,T]: per:state or provinces of residence
[H,T]+Mask: per:state or provinces of residence
Countrywide Financial Corp co-founder Angelo Mozilo retired amidst scandal and investigation. Gold: org:founded by
Mask: no relation
[H,T]: org:top members/employees
[H,T]+Mask: org:founded by
Lt. Assaf Ramon, the son of Israel’s first astronaut, Col. Ilan Ramon, who died in the space shuttle Columbia disaster in 2003, was killed Sunday when an F16-A plane he was piloting crashed in the hills south of Hebron in the West Bank. Gold: per:children
Mask: per:parents
[H,T]: per:siblings
[H,T]+Mask: per:children
Table 6: Comparison of the TACREV test set predictions from the Mask, [H, T] and [H, T] + Mask models that were initialised using RoBERTa-large, focusing on examples where [H,T] confuses the ground truth with a different relation between entities of the same type.
Sentence Label
Asia Bibi was sentenced to hang in Pakistan’s central province of Punjab earlier this month after being accused of insulting the Prophet Mohammed in 2009. Gold: no relation
Token predictions: strained, complicated, tense, fraught, complex.
[H,T]+Mask: no relation
Benjamin Chertoff is the Editor in Chief of Popular Mechanics magazine as well as the cousin of the Director of Homeland Security, Michael Chertoff. Gold: no relation
Token predictions: unclear, unknown, complicated, complex, clear.
[H,T]+Mask: no relation
“The current attempt to restore the commission was masterminded by a suspected mobster, Matteo Messina Denaro, who is among a handful of people vying to replace Provenzano," police said. Gold: per:title
Token predictions: unclear, unknown, known, murky, complex.
[H,T]+Mask: per:title
A man who shot and killed three women in a Pennsylvania health club, then himself, apparently blogged his rage-filled preparations –with the final chilling entry announcing the “big day" of the massacre. Gold: per:state or province of death
Token predictions: unclear, unknown, clear, murky, chilling.
[H,T]+Mask: per:state or province of death
Her brother-in-law, Wen, served as a top Chongqing police official for 16 years before taking over the city’s judiciary. Gold: per:other family
Token predictions: unclear, complicated, unknown, strained, complex.
[H,T]+Mask: per:other family
Robert Holden, deputy director at the National Congress of American Indians, said “the Washington DC-based group is hopeful that the use of secured cards could be expanded to allow tribal members to travel abroad." Gold: org:state or province of headquarters
Token predictions: unclear, confidential, unknown, complex, complicated.
[H,T]+Mask: per:state or provinces of residence
Table 7: Comparison of the tokens predicted for the [MASK] position when using the pre-trained [H, T] + Mask encoder that was initialised using RoBERTa-large. Examples were selected from the TACREV test set. The token predictions are arranged in decreasing order of confidence (score).
Appendix C Ablations and Additional Experiments
Perfomance of the [CLS] token
The intuition behind the Mask strategy is that the embedding of the [MASK] token captures the relational context. Another possible approach is to use the [CLS] token for this purpose. Table 8 analyses a number of variants that are based on this idea. Specifically, we consider the [CLS] token on its own, as well as variants of the [H,T,Mask] and [H,T]+Mask strategies where the role of the [MASK] token is replaced by the [CLS] token. These latter two strategies are referred to as [H,T,CLS] and [H,T]+CLS. For this analysis, we have used TACREV with standard pre-training. Overall, we can draw the same conclusions as for the variants with the [MASK] token, in particular when it comes to comparing [H,T,CLS] with [H,T]+CLS. However, the [CLS] variants yield results which are somewhat worse than the counterparts based on [MASK] (with the exception of the case where [CLS] is used on its own). This justifies the use of a prompt with [MASK].
Impact of Masking Approaches
Peng et al. (2020) highlighted the importance of masking entity spans during pre-training, with some probability, to prevent the model from relying too much on the entity names themselves. Indeed, models which rely too much on the entity names are prone to learning shortcuts which hamper generalisation, a problem which is sometimes referred to as entity bias Wang et al. (2022b). Inspired by these works, we further investigate the impact of different masking strategies during pre-training. Specifically, we look at the following four scenarios in Table 9. First, in the strategy labelled No masking, we do not mask any tokens in the input corpus. Second, in the case of Mask entity spans, for each entity, we mask the entire entity span with 15% probability. For the variant labelled Mask entity span heads, we merely mask out the syntactic heads of entities, for 15% of the entity spans. We find the head word using the SpaCy dependency parser [^6] https://spacy.io/api/dependencyparser. This is motivated by the idea that the syntactic head is most likely to reveal the entity type. Masking this head was found to by beneficial when pre-training entity encoders for this reason Mtumbuka and Schockaert (2023). Finally, in the case of Random tokens, we randomly mask 15% of the tokens in the training corpus. These tokens include both tokens from entity spans and non-entity tokens. This is the strategy that we have used for the main experiments. We can see that masking random tokens gives us the best results.
Comparison of Language Models
In Table 10, we compare the performance of roberta-large, which we have been using for our main experiments, with bert-base-uncased [^7] https://huggingface.co/docs/transformers/model_doc/bert, bert-large-uncased and albert-xxlarge-v1888https://huggingface.co/docs/transformers/model_doc/albert. Unsurprisingly, we can see that bert-large-uncased outperforms bert-base-uncased. Furthermore, across all datasets, roberta-large outperforms bert-large-uncased, but the best results are obtained by albert-xxlarge-v1. This is consistent with the findings from Zhong and Chen (2021). The main advantage of albert-xxlarge-v1 is that it uses 4096-dimensional embeddings, compared to 1024-dimensional embeddings for bert-large-uncased and roberta-large, despite being smaller than the latter two models (due to parameter sharing across layers). As can be seen, for all language models, the main conclusions remain consistent with those reported in the main paper. For instance, we consistently find that the Mask strategy in isolation yields the lowest performance, that [H,T,Mask] improves on [H,T] and that [H,T]+Mask achieves the best results. This further supports our main findings about the complementarity of the [H,T] and Mask representations.
Few-shot Relation Classification
To complement our main results, we have carried out an evaluation in the few-shot setting, where only a few training examples per relation are available. We employ three commonly-used few-shot learning settings on the family of TACRED datasets: 4 training examples per relation (representing approximately 1% of the full TACRED training set), 16 training examples per relation (approximately 5%), and 32 examples per relation (approximately 10%) Sainz et al. (2021); Lu et al. (2022). We compare our models with the methods from Sainz et al. (2021) and Lu et al. (2022), which were specifically designed for the few-shot setting. In particular, Sainz et al. (2021) relied on pre-trained NLI models to solve relation extraction in the few-shot setting. Note that their approach relies on manually constructed verbalisations of the relations, which essentially provides an additional supervision signal. They reported results for NLI models based on RoBERTa-large (shown as $\text{NLI}_{\text{RoBERTa}}$) on DeBERTa-v2xxlarge (shown as $\text{NLI}_{\text{DeBERTa}}$). Lu et al. (2022) treat few-shot relation extraction as a summarisation task, relying on a pretrained PEGASUS-large abstractive summarisation model (shown as $\text{SURE}_{\text{PEGASUS}}$). They also rely on manually constructed verbalisations.
As can be seen in Table 11, the [H,T]+Mask model consistently outperforms $\text{SURE}_{\text{PEGASUS}}$. Furthermore, [H,T]+Mask outperforms $\text{NLI}_{\text{RoBERTa}}$ in all cases apart from the 1% setting, and $\text{NLI}_{\text{DeBERTa}}$ for the 10% and 100% configurations. This is remarkable, given that these baselines were specifically designed for the few-shot setting, rely on extensive pre-training, and in the case of $\text{NLI}_{\text{DeBERTa}}$ and $\text{SURE}_{\text{PEGASUS}}$ rely on much larger LMs. When it comes to the relative performance of the different variants that are considered in this paper, our overall findings are similar as for the main experiments. For instance, [H,T]+Mask and Encore+Mask achieve the best results, the performance of EnCore+Mask is again similar to that of [H,T]+Mask, and pre-training on Gigaword consistently improves the results. The performance of Mask and [H,T] is particularly poor in the few-shot setting, and combining these two types of representations has a very big impact here. For instance, in the setting with 1% of the training data with standard pre-training, the performance increases from 19.7% for [H,T] to 48.4% for [H,T,Mask] and 53.0% for [H,T]+Mask.
TACREV
Mask 22.9
[H, T, Mask] 73.8
[H, T] + Mask 84.6
CLS 25.4
[H, T, CLS] 71.6
[H, T] + CLS 79.2
Table 8: Evaluation of strategies using the [CLS] token on TACREV. For this analysis, we have used RoBERTa-large with the standard pre-training strategy.
TACREV
No masking 58.7
Mask entity spans 73.8
Mask entity span heads 79.4
Mask random tokens 84.6
Table 9: Evaluation of different masking strategies. For this analysis, we have used the [H, T] + Mask approach with RoBERTa-large and standard pre-training.